# Can Large Language Models Act as Ensembler for Multi-GNNs?

### v2

@@ Abstract
Hanqi Duan1, Yao Cheng1, Jianxiang Yu1, Xiang Li1,* 1East China Normal University, Correspondence: xiangli@dase.ecnu.edu.cn (Xiang Li*)
@@ 1 INTRODUCTION
In this paper, we study whether LLMs can act as an ensembler for multi-GNNs and propose the LensGNN model. Given multiple GNNs, the model integrates GNNs with LLMs, avoiding the tedious efforts in cherry-picking GNN across datasets. By dynamically integrating the outputs of different GNNs, LensGNN can not only preserve the semantic richness of node attributes, but also optimize the usage of diverse GNNs, which enhances the model’s generalizability across a wide range of tasks and datasets. Specifically, since multi-GNNs could generate node embeddings in different low-dimensional spaces, we first align multi-GNNs by alternatively feeding their generated representations into a shared classifier to train them sequentially. After that, we freeze the parameters of GNNs to generate node embeddings and integrate them into the embedding layer of the used LLM, treating them as embeddings corresponding to graph tokens that encapsulate graph structural information. Subsequently, we concatenate graph tokens from multi-GNNs, text tokens from node attributes, and hand-crafted instructions to create a prompt and fine-tune the LLM using LoRA. In this way, we can not only align GNNs and LLMs, but implicitly leverage the power of LLM to ensemble multi-GNNs. Finally, the main contributions of our paper are given as follows:
@@ 2 RELATED WORK
Recently, LLMs have revolutionized tasks involving semantic understanding and language generation. Meanwhile, the integration of LLMs and GNNs has garnered significant attention, leading to innovative methodologies Jin et al. (2023). One primary category of approaches takes LLMs as predictors in graph-based tasks, like GraphGPT Tang et al. (2024), LLaGA Chen et al. (2024) GraphLLM Chai et al. (2023) and ENGINE Zhu et al. (2024). Another notable category is to employ LLMs as encoders, as seen in the works of OFA Liu et al. (2024), TextGNN Zhu et al. (2021) and AdsGNN Li et al. (2021), which focus on optimization strategies for encoding graph data. Additionally, the alignment of LLMs and GNNs has been explored through prediction alignment methods, such as GOFA Kong et al. (2024), LTRN Zhang et al. (2021) and GLEM Zhao et al. (2022). These diverse approaches highlight the evolving landscape of LLM and GNN integration. However, there still lack studies exploring multi-GNNs ensembling with LLMs, which is particularly useful to avoid cherry-picking GNNs for a given dataset.
@@ 4.1 Aligning multi-GNNs
$\tilde{h}^{[k]}_{i}=\text{Linear}^{[k]}\left(h^{[k]}_{i}\mid\Theta_{L_{k}}\right).\$ (4)
@@
Multi-GNN alignment. To align node representations from multiple GNNs, we next feed them into a shared linear layer, which serves as the classifier. During the training time, we alternatively input representations from different GNNs into the classifier and train them sequentially. This training approach allows node representations from multi-GNNs to be aligned, which facilitates the integration of multi-GNNs. After training, the parameters of GNNs and the linear layer in Eq. 4 will be frozen, which are then used in aligning GNNs and LLMs. The overall procedure for GNN alignment is summarized in the top of Fig. 2.
In summary, during the multi-GNN alignment step, we train each GNN to ensure that their outputs reside in the same vector space. By incorporating the node feature initialization step that utilizes language pre-trained models, our model gains the ability to extract semantic and structural information from the nodes in the text graph. Building on the full potential of GNNs, we map the dimensions of the GNN representations to the dimensions required by the LLM embeddings, thereby maximizing the integration of extensive graph structural information into the LLM in the subsequent steps.
@@ 4.2 Ensembling multi-GNNs with LLM
[…]
Prompt design and Multi-GNN ensembling. After aligning graph structure and semantics, LLM can understand graph representations. We next design prompts and take LLM as the ensembler to integrate multi-GNNs. Our goal is to leverage the strengths of both multi-GNNs and LLM to output more accurate predictions. The prompt design follows three key principles:
  1. 1.
  2. 2.
  3. 3.
Based on the above principles, we provide a prompt template used in our experiments. Details on the prompt template are given in Appendix D. Further, the two steps: aligning GNNs and LLMs, and ensembing multi-GNNs can be trained simultaneously. In our experiments, we merge them into one step.
@@ 5 EXPERIMENT
Table 1: Comparison on classification accuracy (%) with MLP, GNN models and LM-based methods. LensGNN utilizes Baichuan2-13B-chat as the backbone LLM. We highlight the best score on each dataset in bold and underline the runner-up’s.
@@
LLM Ensembler LensGNN-[GCN+GAT] 88.56 94.11 75.78 78.05 18.41
@@ 5.1 Experimental settings
Baselines. We compare our method with 19 baselines: MLP, GNN models, LM-based models and LLM-based models. Specifically, GNN models include GCN Kipf and Welling (2016), GAT Veličković et al. (2017), GIN Xu et al. (2019), GraphSAGE Hamilton et al. (2017a) and Graphormer Ying et al. (2021). LM-based models consist of BERT Devlin (2018), SentenceBERT Reimers (2019) and DeBERTa He et al. (2020). LLM-based models include DGTL Qin et al. (2024), SNS-GPT4 Li et al. (2024), GAugLLM Fang et al. (2024) and Baichuan2-13B Yang et al. (2023), which all employ 13B LLMs. We also compare LensGNN with other 7B-LLM-based models: GraphGPT Tang et al. (2024), LLaGA Chen et al. (2024), OFA Liu et al. (2024) and GOFA Kong et al. (2024). Due to the space limitation, details on these baselines are given in Appendix B.
Setup. We use Baichuan2 Yang et al. (2023) and InternLM2.5 Cai et al. (2024) as the backbone LLM for LensGNN. In the experiments, we ensemble three widely adopted GNN models: GCN, GAT and GIN, each with two layers. For GNN models, we use node representations obtained from the pre-trained SentenceBERT Reimers (2019) as input. For more details on experiment setup, refer to Appendix C.
@@ 5.2 Node classification results
Since the large number of baselines report their results on different datasets and settings, we show the classification results in three tables, respectively. Table 1 compares LensGNN with MLP, GNN models and small language models. Tables 2 summarizes the results against LLM-based models. Table 3 evaluates the performance of LensGNN in the 20-shot setting. From these tables, we observe:
@@
(2) LensGNN achieves better performance than LLM-based models, even in the 20-shot setting. In particular, we see that Baichuan2-13B performs very poorly. We speculate this is because LLMs that have not been fine-tuned are not suitable for node classification tasks on graph data. Different from other LLM-based models, LensGNN integrates the strengths of multiple GNNs and employs two-phases of alignment, which explains its superior performance.
Table 2: Comparison on classification accuracy (%) with LLM-based models. * denotes the best variant of our model and - means that the results are missing from their original papers.
@@
Table 3: Classification accuracy in 20-shot setting.
@@ 5.3 Graph classification results
Table 4: Graph classification results.
Model BACE BBBP ClinTox
@@
We further evaluate the effectiveness of LensGNN on the graph classification task. We use classification accuracy (ACC) and AUC scores as the evaluation metrics. The results are given in Table  4. From the table, LensGNN outperforms GNN baselines in most cases, which shows that combining GNNs and LLMs is inspiring. Further, LensGNN-ALL generally performs better than its variants. This verifies that leveraging strengths of multiple GNNs is useful. Surprisingly, we find that on BACE, LensGNN and its variants achieve smaller AUC scores than GNN models. We step into the datasets and observe that BACE is a label-balanced dataset while BBBP and ClinTox are highly imbalanced. Note that, for balanced dataset, ACC is more convincing than AUC score, while AUC score is a better indicator than ACC on imbalanced dataset. Therefore, the ACC advantage of LensGNN on BACE and the AUC leads on other two datasets have evidently demonstrated the superiority of our method.
@@ 5.5 Performance with different backbones
(1) Small LMs BERT-base, T5-base, and T5-base (Encoder only) perform well when applied to a single GNN, but the performance drops significantly when ensembling multiple GNN models. This suggests that models with less parameters lack the capability to integrate multiple GNNs.
@@ 6 CONCLUSION
We studied the problem of how to ensemble multi-GNNs in this paper and proposed LensGNN, which ensembles multi-GNNs with LLMs. LensGNN adopts two phases of alignment: the multi-GNN alignment aims to map node representations from different GNNs into the same low-dimensional space, while the GNN-LLM alignment injects graph representation as graph tokens into LLM. After that, we perform supervised fine-tuning with LoRA to enhance the LLM’s capability in understanding graph topology. Finally, we conducted extensive experiments to show the effectiveness of our ensembling model LensGNN. In particular, our experimental results showed that LLMs can serve as effective multi-GNN ensembler, while small LMs cannot.
@@ Appendix B Baselines
(3) LM Based Models: BERT Devlin (2018) is a groundbreaking model in natural language processing. It utilizes the transformer architecture to understand the context of words in a sentence by looking at both their left and right contexts simultaneously. This bidirectional approach enables BERT to capture nuanced meanings and relationships in text. SentenceBERT Reimers (2019) is a modification of the original BERT (Bidirectional Encoder Representations from Transformers) model designed specifically for generating sentence embeddings. It was introduced to improve the performance of BERT on tasks that require understanding the semantic meaning of entire sentences, such as semantic textual similarity, paraphrase identification, and clustering. DeBERTa He et al. (2020) is a transformer-based language model that improves upon BERT and other models like RoBERTa by introducing several innovations.
@@
(4) LLM Based Models: GraphGPT Tang et al. (2024) is a novel framework that integrates Large Language Models (LLMs) with graph structural knowledge through graph instruction tuning. This innovative approach enables LLMs to comprehend and interpret the structural components of graphs, thereby demonstrating superior generalization in both supervised and zero-shot graph learning tasks. LLaGA Chen et al. (2024) integrates the capabilities of LLMs with graph-structured data, enabling LLMs to handle complex graph tasks effectively. LLaGA achieves this by reorganizing graph nodes into structure-aware sequences and mapping them into the token embedding space through a versatile projector, demonstrating superior versatility, generalizability, and interpretability across various datasets and tasks. InstructGLM Ye et al. (2024) use natural language to describe multi-scale geometric structure of the graph and then instruction finetune an LLM to perform graph tasks, which enables Generative Graph Learning. DGTL Qin et al. (2024) incorporates graph structure information through tailored disentangled graph neural network layers, enabling LLMs to capture the intricate relationships hidden in TAGs from multiple structural factors, which is able to enhance the reasoning and predicting capabilities of LLMs for TAGs. SNS-GPT4 Li et al. (2024) is a specialized version of the GPT-4 model designed for Social Network Services (SNS) applications. It introduces similarity-based neighbor selection to effectively improve the quality of selected neighbors, thereby enhancing graph representation and alleviating issues like over-squashing and heterophily. GAugLLM Fang et al. (2024) is a novel framework for augmenting TAGs. It leverages advanced large language models like Mistral to enhance self-supervised graph learning. Baichuan2 Yang et al. (2023) refers to the second iteration of the Baichuan large language model (LLM). It has been trained on an impressive 2.6 trillion high-quality tokens, ensuring a robust understanding of language nuances. Baichuan 2 comes in two main versions: 7B and 13B, both available in Base and Chat configurations, with the latter offering 4-bit quantization for efficient deployment. OFA Liu et al. (2024) is a pioneering framework that unifies various graph tasks into a single model, enabling it to address classification tasks across different domains and tasks. OFA achieves this by converting graph data into text-attributed graphs (TAGs), using language models to encode diverse text attributes into a common embedding space, and introducing a novel graph prompting paradigm for in-context learning without fine-tuning. GOFA Kong et al. (2024) is a novel graph foundation model that integrates the strengths of large language models (LLMs) and graph neural networks (GNNs) to enable joint graph and language modeling. It achieves this by interleaving GNN layers into a frozen pre-trained LLM, allowing for self-supervised pretraining on graph data and fluidity in handling various graph-related tasks.
Table 8: Hyperparameter Search Spaces.

### v3

@@ Abstract
Hanqi Duan1, Yao Cheng1, Jianxiang Yu1, Yao Liu1, Xiang Li1,* 1East China Normal University, Correspondence: xiangli@dase.ecnu.edu.cn (Xiang Li*)
@@ 1 INTRODUCTION
In this paper, we study whether LLMs can act as an ensembler for multi-GNNs and propose the LensGNN model. Given multiple GNNs, the model ensembles GNNs with LLMs, avoiding the tedious efforts in cherry-picking GNN across datasets. By dynamically integrating the outputs of different GNNs, LensGNN can not only preserve the semantic richness of node attributes, but also optimize the usage of diverse GNNs, which enhances the model’s generalizability across a wide range of tasks and datasets. Specifically, since multi-GNNs could generate node embeddings in different low-dimensional spaces, we first align multi-GNNs by alternatively feeding their generated representations into a shared classifier to train them sequentially. After that, we freeze the parameters of GNNs to generate node embeddings and integrate them into the embedding layer of the used LLM, treating them as embeddings corresponding to graph tokens that encapsulate graph structural information. Subsequently, we concatenate graph tokens from multi-GNNs, text tokens from node attributes, and hand-crafted instructions to create a prompt and fine-tune the LLM using LoRA. In this way, we can not only align GNNs and LLMs, but implicitly leverage the power of LLM to ensemble multi-GNNs. Finally, the main contributions of our paper are given as follows:
@@ 2 RELATED WORK
Recently, LLMs have revolutionized tasks involving semantic understanding and language generation. Meanwhile, the ensemble of LLMs and GNNs has garnered significant attention, leading to innovative methodologies Jin et al. (2023). One primary category of approaches takes LLMs as predictors in graph-based tasks, like GraphGPT Tang et al. (2024), LLaGA Chen et al. (2024) GraphLLM Chai et al. (2023) and ENGINE Zhu et al. (2024). Another notable category is to employ LLMs as encoders, as seen in the works of OFA Liu et al. (2024), TextGNN Zhu et al. (2021) and AdsGNN Li et al. (2021), which focus on optimization strategies for encoding graph data. Additionally, the alignment of LLMs and GNNs has been explored through prediction alignment methods, such as GOFA Kong et al. (2024), LTRN Zhang et al. (2021) and GLEM Zhao et al. (2022). These diverse approaches highlight the evolving landscape of LLM and GNN ensemble. However, there still lack studies exploring multi-GNNs ensembling with LLMs, which is particularly useful to avoid cherry-picking GNNs for a given dataset.
@@ 4.1 Aligning multi-GNNs
$\tilde{h}^{[k]}_{i}=\text{Linear}^{[k]}\left(h^{[k]}_{i}\mid\Theta_{L_{k}}\right).\\$ (4)
@@
Multi-GNN alignment. To align node representations from multiple GNNs, we next feed them into a shared MLP, which serves as the classifier. During the training time, we alternatively input representations from different GNNs into the classifier and train them sequentially. This training approach allows node representations from multi-GNNs to be aligned, which facilitates the ensemble of multi-GNNs. For better alignment, we freeze this MLP starting from a certain training epoch. After training, the parameters of GNNs and the linear layer in Eq. 4 will be frozen, which are then used in aligning GNNs and LLMs. The overall procedure for GNN alignment is summarized in the top of Fig. 2.
In summary, during the multi-GNN alignment step, we train each GNN to ensure that their outputs reside in the same vector space. By incorporating the node feature initialization step that utilizes language pre-trained models, our model gains the ability to extract semantic and structural information from the nodes in the text graph. Building on the full potential of GNNs, we map the dimensions of the GNN representations to the dimensions required by the LLM embeddings, thereby maximizing the ensemble of extensive graph structural information into the LLM in the subsequent steps.
@@ 4.2 Ensembling multi-GNNs with LLM
To address the problem, recent advancements in LoRA (Low-Rank Adaptation) Hu et al. (2021) training have been proposed to directly fine-tune LLMs for alignment. LoRA is an efficient fine-tuning technique for large pre-trained models, and introduces low-rank matrices to simulate the effects of full parameter tuning, significantly reducing the number of training parameters and computational resources required. This allows for effective customization of LLMs even with limited resources. Thus, we adopt LoRA training to fine-tune the LLMs, enabling them to comprehend GNN representations. Specifically, given an instruction prompt, we reserve several dummy tokens in the prompt text at the appropriate positions where the GNN representations need to be inserted. For example, in the prompt ‘(GNN type: GCN, representations: $\text{graph\_token}_{1}>,<\text{graph\_token}_{2}>,\ldots,<\text{graph\_token}_{t}>$ ), which category does it belong to?’, $\text{graph\_token}_{i}>$ denotes the $i$-th dummy token. Then, for each token in the prompt text, LLM can generate an embedding vector of size $e$ . These vectors can form an embedding for the whole prompt text in the format of $\ldots,E^{\prime}_{1},E^{\prime}_{2},\ldots,E^{\prime}_{t},\ldots]$, where $E^{\prime}_{i}$ is the embedding vector of $\text{graph\_token>}_{i}$. After that, we overwrite the embeddings of dummy tokens by GNN representations. From Eq. 4, we can generate a GNN representation vector of length $t\times e$, which can be further segmented into $t$ one-dimensional vectors of length $e$, denoted as $E^{G}_{1},E^{G}_{2},\ldots,E^{G}_{t}]$. These vectors serve as the embeddings of $t$ graph tokens. After substitution, the raw embedding vector of the prompt text becomes $\ldots,E^{G}_{1},E^{G}_{2},\ldots,E^{G}_{t},\ldots]$. We next feed it into LLM and use LoRA for fine-tuning. In this way, we can automatically leverage the language understanding capability of LLM, and implicitly align structure and semantics.
Prompt design and Multi-GNN ensembling. After aligning graph structure and semantics, LLM can understand graph representations. We next design prompts and take LLM as the ensembler to integrate multi-GNNs. Our goal is to leverage the strengths of both multi-GNNs and LLM to output more accurate predictions.
The designed prompts include three major components: task instruction, node text and graph token. Specifically, task instruction consists of task specifications and also instructions that guide LLMs to ensemble multi-GNNs. Node text is used to inject textual attributes of target nodes and their neighbors. Due to the input length limitation, for each target node, we only consider up to 20 adjacent nodes. When the number of adjacent nodes for the target node exceeds 20 or when the total token count of the prompt surpasses the predefined limit, a random sampling of the adjacent nodes is performed. Further, graph tokens are derived from multi-GNNs, which contain rich graph structural information. After prompts are deigned, they are fed into LLMs to predict labels. Formally, we have:
$o=LLM(x_{\text{instruction}},x_{\text{node-text}},\{x_{\text{GNN}_{i}}\}_{i=1}^{k}),$
where $o$ is the predicted label, $x_{\text{instruction}}$ denotes the instruction text, $x_{\text{node-text}}$ is node textual descriptions and $x_{\text{GNN}_{i}}$ denotes the graph representation (graph token) generated by the $i$-th GNN. Details on the prompt template used in our experiments are given in Appendix D. Therefore, the two steps: aligning GNNs and LLMs, and ensembing multi-GNNs can be trained simultaneously. In our experiments, we merge them into one step.
@@ 5 EXPERIMENT
SciBERT 83.21 95.26 73.11 77.74 76.83
MedBERT 77.31 93.94 72.09 74.29 74.04
LLM Ensembler
LensGNN-[GCN+GAT] 88.56 94.11 75.78 78.05 18.41
@@
Table 1: Comparison on classification accuracy (%) with MLP, GNN models and LM-based methods. LensGNN utilizes Baichuan2-13B-chat as the backbone LLM. We highlight the best score on each dataset in bold and underline the runner-up’s. Here, LensGNN-ALL refers to LensGNN-[GCN+GAT+GIN].
@@ 5.1 Experimental settings
Baselines. We compare our method with 23 baselines: MLP, GNN models, ensemble models, LM-based models and LLM-based models. Specifically, GNN models include GCN, GAT, GIN, GraphSAGE Hamilton et al. (2017a) and Graphormer Ying et al. (2021). Ensemble models include Bagging Breiman (1996), Stacking Wolpert (1992) and AdaBoost Freund and Schapire (1997). LM-based models consist of BERT Devlin (2018), SentenceBERT Reimers (2019), DeBERTa He et al. (2020), SciBERT Beltagy et al. (2019) and MedBERT Feihong Yang (2021). LLM-based models include DGTL Qin et al. (2024), SNS-GPT4 Li et al. (2024), GAugLLM Fang et al. (2024) and Baichuan2-13B Yang et al. (2023), which all employ 13B LLMs. We also compare LensGNN with other 7B-LLM-based models: GraphGPT Tang et al. (2024), LLaGA Chen et al. (2024), InstructGLM Ye et al. (2024), OFA Liu et al. (2024) and GOFA Kong et al. (2024). Due to the space limitation, details on these baselines are given in Appendix B.
Setup. We use Baichuan2 and InternLM2.5 Cai et al. (2024) as the backbone LLM for LensGNN. In the experiments, we ensemble three widely adopted GNN models: GCN, GAT and GIN, each with two layers. For GNN models, we use node representations obtained from the pre-trained SentenceBERT Reimers (2019) as input. The ensemble experiments for more GNNs are presented in Appendix I. The hyperparameter analysis experiment is detailed in Appendix J. The analysis of input length is presented in Appendix K. For more details on experiment setup, refer to Appendix C. Detailed hyperparameter settings are provided in Appendix L. Additionally, in Appendix H, we include supplementary experiments utilizing LLama2-7B as the backbone.
@@ 5.2 Node classification results
Since the large number of baselines report their results on different datasets and settings, we show the classification results in three tables, respectively. Table 1 compares LensGNN with MLP, GNN models and small language models. Tables 2 summarizes the results against LLM-based models. Table 3 evaluates the performance of LensGNN in the 20-shot setting. Table 4 provides a comparison between LensGNN and methods that ensemble GCN, GAT and GIN. We also experimented with traditional methods that ensemble LLM and GNNs. The experimental analysis is provided in Appendix G. Additionally, we conduct one-shot experiments, which are discussed in Appendix F. From these tables, we observe:
@@
(2) LensGNN achieves better performance than LLM-based models, even in the 20-shot setting. In particular, we see that Baichuan2-13B performs very poorly. We speculate this is because LLMs that have not been fine-tuned are not suitable for node classification tasks on graph data. Different from other LLM-based models, LensGNN ensembles the strengths of multiple GNNs and employs two-phases of alignment, which explains its superior performance.
(3) LensGNN leads standard ensemble methods across all datasets. This is because when ensembling multiple GNNs, standard ensemble methods can only leverage graph structural information. However, LLMs as ensembler have strong semantic understanding capability that further improves the model performance.
@@
Table 2: Comparison on classification accuracy (%) with LLM-based models. * denotes the best variant of our model and - means that the results are missing from their original papers. The best variant model is given in Appendix E.
@@
Table 3: Classification accuracy (%) in 20-shot setting.
Model Cora PubMed Citeseer WikiCS
Bagging 87.16 82.67 77.35 76.97
Stacking 86.25 84.85 78.08 74.12
AdaBoost 84.91 83.36 76.09 75.41
LensGNN-ALL 91.88 95.68 79.31 83.47
Table 4: Comparison with ensemble methods (%).
@@ 5.3 Graph classification results
Model
BACE BBBP ClinTox
@@
Table 5: Graph classification results.
We further evaluate the effectiveness of LensGNN on the graph classification task. We use classification accuracy (ACC) and AUC scores as the evaluation metrics. The results are given in Table  5. From the table, LensGNN outperforms GNN baselines in most cases, which shows that combining GNNs and LLMs is inspiring. Further, LensGNN-ALL generally performs better than its variants. This verifies that leveraging strengths of multiple GNNs is useful. Surprisingly, we find that on BACE, LensGNN and its variants achieve smaller AUC scores than GNN models. We step into the datasets and observe that BACE is a label-balanced dataset while BBBP and ClinTox are highly imbalanced. Note that, for balanced dataset, ACC is more convincing than AUC score, while AUC score is a better indicator than ACC on imbalanced dataset. Therefore, the ACC advantage of LensGNN on BACE and the AUC leads on other two datasets have evidently demonstrated the superiority of our method.
@@ 5.4 Ablation study
On Cora and Wiki-CS, the small training sets introduce more noise during training. Since our model relies on an LLM as backbone—and LLMs typically require large-scale data for effective learning—we attribute the underperformance of LensGNN-ALL compared to LensGNN-[GCN+GIN] primarily to the limited training data and elevated noise levels. These factors hinder the LLM from adequately learning to align all three GNNs (GCN, GAT, and GIN) simultaneously.
@@ 5.5 Performance with different backbones
(1) Small LMs BERT-base, T5-base, and T5-base (Encoder only) perform well when applied to a single GNN, but the performance drops significantly when ensembling multiple GNN models. This suggests that models with less parameters lack the capability to ensemble multiple GNNs.
@@ 6 CONCLUSION
We studied the problem of how to ensemble multi-GNNs in this paper and proposed LensGNN, which ensembles multi-GNNs with LLMs. LensGNN adopts two phases of alignment: the multi-GNN alignment aims to map node representations from different GNNs into the same low-dimensional space, while the GNN-LLM alignment injects graph representation as graph tokens into LLM. After that, we performed supervised fine-tuning with LoRA to enhance the LLM’s capability in understanding graph topology. Finally, we conducted extensive experiments to show the effectiveness of our ensembling model LensGNN. In particular, the results show that LLMs can serve as effective multi-GNN ensembler, while small LMs cannot.
@@ Appendix B Baselines
(3) Ensemble Models: Bagging Breiman (1996) is an ensemble method that improves model stability and accuracy by training multiple base learners on bootstrap samples and aggregating their predictions through voting or averaging. Stacking Wolpert (1992) is an ensemble technique that combines predictions from multiple base learners to train a meta-learner, enhancing overall predictive performance through stacked generalization. AdaBoost Freund and Schapire (1997) is an adaptive boosting algorithm that sequentially trains weak learners, adjusts sample weights, and combines their predictions to create a strong classifier with improved accuracy.
(4) LM Based Models: BERT Devlin (2018) is a groundbreaking model in natural language processing. It utilizes the transformer architecture to understand the context of words in a sentence by looking at both their left and right contexts simultaneously. This bidirectional approach enables BERT to capture nuanced meanings and relationships in text. SentenceBERT Reimers (2019) is a modification of the original BERT (Bidirectional Encoder Representations from Transformers) model designed specifically for generating sentence embeddings. It was introduced to improve the performance of BERT on tasks that require understanding the semantic meaning of entire sentences, such as semantic textual similarity, paraphrase identification, and clustering. DeBERTa He et al. (2020) is a transformer-based language model that improves upon BERT and other models like RoBERTa by introducing several innovations. SciBERT Beltagy et al. (2019) is a BERT-based language model pre-trained on a large multi-domain corpus of scientific publications to achieve state-of-the-art performance on various scientific NLP tasks. MedBERT Feihong Yang (2021) is a series of BERT and ALBERT models pre-trained on a large Chinese clinical corpus, which outperforms baseline models on named entity recognition and text classification tasks within the Chinese clinical NLP domain.
@@
(5) LLM Based Models: GraphGPT Tang et al. (2024) is a novel framework that integrates Large Language Models (LLMs) with graph structural knowledge through graph instruction tuning. This innovative approach enables LLMs to comprehend and interpret the structural components of graphs, thereby demonstrating superior generalization in both supervised and zero-shot graph learning tasks. Additionally, GraphGPT employs Baichuan-7B and Vicuna-7B as its backbone models. LLaGA Chen et al. (2024) integrates the capabilities of LLMs with graph-structured data, enabling LLMs to handle complex graph tasks effectively. LLaGA achieves this by reorganizing graph nodes into structure-aware sequences and mapping them into the token embedding space through a versatile projector, demonstrating superior versatility, generalizability, and interpretability across various datasets and tasks. LLaGA employs Vicuna-7B as its backbone model. InstructGLM Ye et al. (2024) use natural language to describe multi-scale geometric structure of the graph and then instruction finetune an LLM to perform graph tasks, which enables Generative Graph Learning. InstructGLM employs Llama2-7B as its backbone model. DGTL Qin et al. (2024) is an LLM-first-GNN-later method, which means first using LLMs to process text before training GNNs. DGTL incorporates graph structure information through tailored disentangled graph neural network layers, enabling LLMs to capture the intricate relationships hidden in TAGs from multiple structural factors, which is able to enhance the reasoning and predicting capabilities of LLMs for TAGs. DGTL employs Llama2-13B as its backbone model. SNS-GPT4 Li et al. (2024) is a specialized version of the GPT-4 model designed for Social Network Services (SNS) applications. It introduces similarity-based neighbor selection to effectively improve the quality of selected neighbors, thereby enhancing graph representation and alleviating issues like over-squashing and heterophily. SNS-GPT4 employs GPT3.5-turbo and GPT4 as its backbone models. Notably, SNS-GPT4 demonstrates better performance than TAPE He et al. (2023) (another LLM-first-GNN-later approach) in its reported results. GAugLLM Fang et al. (2024) is a novel framework for augmenting TAGs. It leverages advanced large language models like Mistral to enhance self-supervised graph learning. GAugLLM employs Mistral 8*7b, ChatGPT-3.5 and Llama2-13b as its backbone models. Baichuan2 Yang et al. (2023) refers to the second iteration of the Baichuan large language model (LLM). It has been trained on an impressive 2.6 trillion high-quality tokens, ensuring a robust understanding of language nuances. Baichuan 2 comes in two main versions: 7B and 13B, both available in Base and Chat configurations, with the latter offering 4-bit quantization for efficient deployment. OFA Liu et al. (2024) is a pioneering framework that unifies various graph tasks into a single model, enabling it to address classification tasks across different domains and tasks. OFA achieves this by converting graph data into text-attributed graphs (TAGs), using language models to encode diverse text attributes into a common embedding space, and introducing a novel graph prompting paradigm for in-context learning without fine-tuning. OFA employs Llama2 as its backbone model. GOFA Kong et al. (2024) is a novel graph foundation model that integrates the strengths of large language models (LLMs) and graph neural networks (GNNs) to enable joint graph and language modeling. It achieves this by interleaving GNN layers into a frozen pre-trained LLM, allowing for self-supervised pretraining on graph data and fluidity in handling various graph-related tasks. OFA employs Llama2 and Mistral as its backbone models.
Appendix C Experiment setup
In addition, in the era of LLMs, due to the high cost, it is difficult to run experiments for baselines. For fair comparison, a widely adopted approach is to make comparison on the results for baselines reported from their original papers because their authors have well fine-tuned these models. This explains why only part of results are included in Table 2 and why we did not reimplement some baselines.